# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ B.2 Motivations: Reductions to and generalisations of existing settings
Fairness of exposure in bandits. Wang et al., 2021a introduced positive merit based exposure of fairness constraints (Singh and Joachims,, 2019) in stochastic bandits standing against the winner-takes-all allocation strategy that are historically studied. The chosen allocation in this setting should satisfy the fairness constraint $\frac{\boldsymbol{\pi}^{\star}_{a}}{f(\mu^{\star}_{a})}=\frac{\boldsymbol{\pi}^{\star}_{a^{\prime}}}{f(\mu^{\star}_{a^{\prime}})},\forall a^{\prime}\in[{K}]$ where $f(.)$ transform reward of an arm to a positive merit. Though Wang et al., 2021a studied this setting in regret analysis, this setting in BAI setting is a direct application of our setting as we are looking for an optimal policy $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}$ such that $\boldsymbol{\pi}^{\star}$ satisfies $\mathbf{A}_{f({\boldsymbol{\mu}})}^{T}\boldsymbol{\pi}=0$ where $\mathbf{A}_{f({\boldsymbol{\mu}})}$ is of order $\frac{K(K-1)}{2}\times K$ and $\mathbf{A}_{f({\boldsymbol{\mu}})}$ is expressed as,
$(\mathbf{A}_{f({\boldsymbol{\mu}})})_{ij}=\begin{cases}\frac{1}{f(\mu_{a})}&\quad\text{if }aK-\frac{1}{2}(a-1)(a-2)\leq j\leq aK-\frac{1}{2}a(a-1)\text{ and }i=a\,,\\ \frac{-1}{f(\mu_{j})}&\quad\text{if }aK-\frac{1}{2}(a-1)(a-2)\leq j\leq aK-\frac{1}{2}a(a-1)\text{ and }a<i\leq K\,,\\ 0&\quad\text{otherwise.}\\ \end{cases}$
For example, when $K=3$, $\mathbf{A}_{f({\boldsymbol{\mu}})}=[[\frac{1}{\mu_{1}},-\frac{1}{\mu_{2}},0],[0,\frac{1}{\mu_{2}},-\frac{1}{\mu_{3}}],[\frac{1}{\mu_{1}},0,-\frac{1}{\mu_{3}}]]$.
@@ Appendix C Strong Duality and the Lagrangian multiplier: Proof of Theorem 1
$\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{\mathbf{A}^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{T}{\mathbf{A}^{\prime}}\boldsymbol{\omega}$
$=$ $\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$
Here, $\mathcal{L}\mathrel{\triangleq}\{\boldsymbol{l}\in\mathbb{R}^{d}\mid 0\leq\|{\boldsymbol{l}}\|_{1}\leq\frac{1}{\gamma}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\}$, where $\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{\mathbf{A}}^{i}\boldsymbol{\omega}^{*}\}$, i.e. the minimum slack for pessimistic constraints w.r.t. the optimal allocation.
@@
This proof involves three steps. In the first step we prove convexity and other properties of the sets involved in the main optimisation problem 1. In the next step, we show that Slater’s sufficient conditions hold for ${\boldsymbol{\pi}}$ as a consequence of these properties. Once we prove the unique optimality of ${\boldsymbol{\pi}}$ we state bounds on the L1-norm of the Lagrangian multiplier. We conclude by establishing strong duality and proving the statement of the theorem.
Step 1: Properties of perturbed feasible set and alt-set. Let us first check the properties of $\hat{\mathcal{F}}$, $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ and $\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})$. For that, let us remind the definitions of these sets. The estimated feasible set is defined as ^$\hat{\mathcal{F}}\mathrel{\triangleq}\left\{{\boldsymbol{\pi}}\in\Delta_{K}:\tilde{\mathbf{A}}{\boldsymbol{\pi}}\leq 0\right\}$. The set of alternative (confusing) instances for the optimal policy $\hat{\boldsymbol{\pi}}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$ is $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})\mathrel{\triangleq}\left\{{\boldsymbol{\lambda}}\in\mathbb{D}:r+\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}{\boldsymbol{\lambda}}^{T}{\boldsymbol{\pi}}>{\boldsymbol{\lambda}}^{T}\hat{\boldsymbol{\pi}}^{\star}\right\}$. For $\boldsymbol{\pi}^{\prime}$ being a neighbour of $\hat{\boldsymbol{\pi}}^{\star}$ or in other words, an extreme point in $\hat{\mathcal{F}}$, we decompose the alternative set as the union of half-spaces as,
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})=\bigcup_{{\boldsymbol{\pi}}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})}\left\{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{T}(\hat{\boldsymbol{\pi}}^{\star}-{\boldsymbol{\pi}}^{\prime})<r\right\}$
We should note, $\boldsymbol{\pi}^{\prime}$ shares at least $({K}-1)$ active constraints with $\hat{\boldsymbol{\pi}}^{\star}$. It is clear that $\hat{\mathcal{F}}$ is bounded and convex in ${\boldsymbol{\pi}}$. Since, convex combination of any two extreme point ${\boldsymbol{\pi}}^{\prime}_{1},{\boldsymbol{\pi}}^{\prime}_{2}$ in the neighborhood of the optimal policy $\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})$ also shares ${K}-1$ active constraints with $\hat{\boldsymbol{\pi}}^{\star}$, so $\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})$ is convex in $\boldsymbol{\pi}^{\prime}$.
@@
Therefore, any convex combination of these neighbourhood policy, ${\boldsymbol{\lambda}}^{T}\left(\hat{\boldsymbol{\pi}}^{\star}-(a{\boldsymbol{\pi}}^{\prime}_{1}+(1-a){\boldsymbol{\pi}}^{\prime}_{2})\right)={\boldsymbol{\lambda}}^{T}(\hat{\boldsymbol{\pi}}^{\star}-{\boldsymbol{\pi}}^{\prime}_{2})-a{\boldsymbol{\lambda}}^{T}({\boldsymbol{\pi}}^{\prime}_{1}-{\boldsymbol{\pi}}^{\prime}_{2})\leq c$. Therefore, the set $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is also bounded and convex in ${\boldsymbol{\pi}}$.
Also, since we are working with pessimistic estimate of $\mathbf{A}$, the set $\hat{\mathcal{F}}$ will always be non-empty, because we will find at least one $\tilde{\mathbf{A}}_{0}$ which is non-singular and it’s inverse exists.
@@
    $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ is also bounded and convex in ${\boldsymbol{\pi}}$.
@@
$0\leq||{l^{\star}}||_{1}\leq\frac{1}{\gamma}\left(\mathcal{D}(\boldsymbol{\omega}^{\star},{\boldsymbol{\mu}},\hat{\mathcal{F}})-\mathcal{D}(\boldsymbol{\omega}^{\star},\hat{{\boldsymbol{\mu}}},\mathcal{F})\right)\leq\frac{1}{\gamma}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$
@@
Where, $\hat{\boldsymbol{\pi}}^{\star}$ is the $r$-good feasible policy.
Step 4: Establishing strong duality. Therefore the domain of the Lagrangian multiplier is also bounded and convex. So again we say that $l_{t}^{\star}$ uniquely minimises Equation 1. We define $\mathcal{L}\mathrel{\triangleq}\{\boldsymbol{l}\in\mathbb{R}^{d}\mid 0\leq\|{\boldsymbol{l}}\|_{1}\leq\frac{1}{\gamma}T_{\hat{\mathcal{F}},r}^{-1}(\hat{{\boldsymbol{\mu}}})\}$, where $\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{\mathbf{A}}^{i}\boldsymbol{\omega}^{*}\}$ Then according to Heine-Borel’s theorem (Theorem 9) we can say that these sets are compact as well. We can then conclude that Strong duality holds which means that it perfectly make sense of solving the Lagrangian dual formulation of the primal optimisation problem because there is no duality gap. We later on will consider this formulation as two player zero sum game. Due to strong duality we claim that the agent wile playing this game, Nash equilibrium will be eventually established.
@@
$\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{\mathbf{A}^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{T}{\mathbf{A}^{\prime}}\boldsymbol{\omega}$
$=$ $\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$
@@ Appendix D Lagrangian Relaxation of Projection Lemma: Proof of Theorem 2
For a sequence {^$\{\hat{\mathcal{F}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{{\boldsymbol{\lambda}}}_{t}\}_{t\in\mathbb{N}}$, we first show that (a) $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, (b) ${\boldsymbol{\lambda}}^{\star}$ is unique, and (c) $\lim_{t\rightarrow\infty}\hat{{\boldsymbol{\lambda}}}_{t}\rightarrow{\boldsymbol{\lambda}}^{\star}$.
Thus, for any $\boldsymbol{\omega}\in\mathcal{F}$ and ${\boldsymbol{\mu}}$,
@@
where ${\boldsymbol{\lambda}}^{*}$ is such that for any ${\boldsymbol{\lambda}}^{*}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$
@@
Statement (a): Convergence of the limit $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}$. To begin with the proof of the first statement of Theorem 2 we leverage the results stated in Theorem 10. Let $H(\tilde{\mathbf{A}})\mathrel{\triangleq}\left\{{\boldsymbol{\pi}}\in\Delta_{K}:\tilde{\mathbf{A}}{\boldsymbol{\pi}}\leq 0\right\}$ and the set function $\tilde{\mathbf{A}}\rightarrow H(\tilde{\mathbf{A}})\cap C$ where $C=\hat{\mathcal{F}}$ is a non-empty compact (proven in Section C subset of $\Delta_{K}$ Then the set $H(\tilde{\mathbf{A}})\cap C$ can be written as
$H(\tilde{\mathbf{A}})\cap C=\left\{{\boldsymbol{\pi}}\in\hat{\mathcal{F}}:\tilde{\mathbf{A}}{\boldsymbol{\pi}}\leq 0\right\}$
To apply Theorem 10, $\{\tilde{\mathbf{A}}^{r},r\in\mathbb{N}\}$, must be a convergent sequence of affine function. It is evident that $\tilde{\mathbf{A}}^{r}$ for any $r\in\mathbb{N}$ is an affine function since $\mathbf{A}$ is linear in $\mathbf{A}$ and the induced pessimism works as a translation. Then we can proceed to the next part of the proof of statement 1 where we prove that $\{\tilde{\mathbf{A}}^{r}\}_{r\in\mathbb{N}}$ is a convergent sequence of functions. For ease of notation we will denote $\tilde{\mathbf{A}}_{t}$ for the t-th element of the sequence $\{\tilde{\mathbf{A}}^{r}\}_{r\in\mathbb{N}}$ for $t\in\mathbb{N}$.
@@
$\mathbb{P}\left(\hat{\mathbf{A}}_{t}^{i}-f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}\leq\mathbf{A}^{i}\leq\hat{\mathbf{A}}_{t}^{i}+f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}\right)\geq 1-\delta$
$\implies$ $\mathbb{P}\left(-\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\leq\frac{\hat{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\leq\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\right)\geq 1-\delta$
$\implies$ $\mathbb{P}\left(-\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\leq Z\leq\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\right)\geq 1-\delta$
$\implies$ $2\Phi\left(\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\right)\geq 2-\delta$
$\implies$ $\frac{f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}\geq\Phi^{-1}\left(1-\frac{\delta}{2}\right)$
$\implies$ $f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}\geq\sigma(\hat{\mathbf{A}}_{t}^{i})\Phi^{-1}\left(1-\frac{\delta}{2}\right)$
where $Z\mathrel{\triangleq}\frac{\hat{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i}}{\sigma(\hat{\mathbf{A}}_{t}^{i})}$ and $\Phi(.)$ is the cumulative distribution function of $\mathcal{N}(0,1)$ distribution. $\lim_{t\rightarrow\infty}f(t,\delta)\|\boldsymbol{\omega}_{t}\|_{\Sigma_{t}^{-1}}\rightarrow 0$ since $\sigma(\hat{\mathbf{A}}_{t}^{i})=\mathcal{O}(\sqrt{\frac{\log t}{t}})$. Leveraging CLT at this point we say
@@
It implies that $\{\tilde{\mathbf{A}}_{r}\}_{r\in\mathbb{N}}$ is a convergent sequence of function for $\mathbf{A}$. Now, we use Theorem 10 and get the following properties of the feasible set.
@@
    $H(\tilde{\mathbf{A}})\cap C\subset\lim_{r\rightarrow\infty}H(\tilde{\mathbf{A}}^{r})\cap C$
@@
    $\lim_{r\rightarrow\infty}H(\tilde{\mathbf{A}}^{r})\cap C$ is a closed convex superset of $H(\tilde{\mathbf{A}})\cap C$.
@@
    $H(\tilde{\mathbf{A}})\cap C$ has non-empty interior because of the feasibility condition and no component in $\tilde{\mathbf{A}}$ is identically $\mathbf{0}$.
    $\lim_{r\rightarrow\infty}H(\tilde{\mathbf{A}}^{r})\cap C=H(\tilde{\mathbf{A}})\cap C$

### v2

@@ B.2 Motivations: Reductions to and Generalisations of The Existing BAI Settings
Fairness of exposure in bandits. Wang et al., 2021a introduced positive merit based exposure of fairness constraints (Singh and Joachims,, 2019) in stochastic bandits standing against the winner-takes-all allocation strategy that are historically studied. The chosen allocation in this setting should satisfy the fairness constraint $\frac{\boldsymbol{\pi}^{\star}_{a}}{f(\mu^{\star}_{a})}=\frac{\boldsymbol{\pi}^{\star}_{a^{\prime}}}{f(\mu^{\star}_{a^{\prime}})},\forall a^{\prime}\in[{K}]$ where $f(.)$ transform reward of an arm to a positive merit. Though Wang et al., 2021a studied this setting in regret analysis, this setting in BAI setting is a direct application of our setting as we are looking for an optimal policy $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$ such that $\boldsymbol{\pi}^{\star}$ satisfies $A_{f({\boldsymbol{\mu}})}^{\top}\boldsymbol{\pi}=0$ where $A_{f({\boldsymbol{\mu}})}$ is of order $\frac{K(K-1)}{2}\times K$ and $A_{f({\boldsymbol{\mu}})}$ is expressed as,
[…]
For example, when $K=3$, $A_{f({\boldsymbol{\mu}})}=[[\frac{1}{\mu_{1}},-\frac{1}{\mu_{2}},0],[0,\frac{1}{\mu_{2}},-\frac{1}{\mu_{3}}],[\frac{1}{\mu_{1}},0,-\frac{1}{\mu_{3}}]]$.
@@ Appendix C Strong Duality and the Lagrangian Multiplier: Proof of Theorem 1
$\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{A^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}{A^{\prime}}\boldsymbol{\omega}$
$=$ $\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\,.$
Here, $\mathcal{L}\mathrel{\triangleq}\{\boldsymbol{l}\in\mathbb{R}^{d}\mid 0\leq\|{\boldsymbol{l}}\|_{1}\leq\frac{1}{\gamma}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})\}$, where $\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{A}^{i}\hat{\boldsymbol{\pi}^{\star}}\}$, i.e. the minimum slack for optimistic constraints w.r.t. the estimated optimal feasible policy.
@@
This proof involves three steps. In the first step we prove convexity and other properties of the sets involved in the main optimisation problem 7. In the next step, we show that Slater’s sufficient conditions hold for $\boldsymbol{\pi}$} as a consequence of these properties. Once we prove the unique optimality of $\boldsymbol{\pi}$} we state bounds on the L1-norm of the Lagrangian multiplier. We conclude by establishing strong duality and proving the statement of the theorem.
[…]
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})=\bigcup_{{\boldsymbol{\pi}}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})}\left\{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-{\boldsymbol{\pi}}^{\prime})<r\right\}$
We should note, $\boldsymbol{\pi}^{\prime}$ shares at least (K1)({K}-1) active constraints with $\hat{\boldsymbol{\pi}}^{\star}$. It is clear that $\hat{\mathcal{F}}$ is bounded and convex in $\boldsymbol{\pi}$}. Since, convex combination of any two extreme point ${\boldsymbol{\pi}}^{\prime}_{1},{\boldsymbol{\pi}}^{\prime}_{2}$ in the neighborhood of the optimal policy $\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})$ also shares K1{K}-1 active constraints with $\hat{\boldsymbol{\pi}}^{\star}$, so $\nu_{\hat{\mathcal{F}}}(\hat{\boldsymbol{\pi}}^{\star})$ is convex in $\boldsymbol{\pi}^{\prime}$.
@@
Therefore, any convex combination of these neighbourhood policy, ${\boldsymbol{\lambda}}^{\top}\left(\hat{\boldsymbol{\pi}}^{\star}-(a{\boldsymbol{\pi}}^{\prime}_{1}+(1-a){\boldsymbol{\pi}}^{\prime}_{2})\right)={\boldsymbol{\lambda}}^{\top}(\hat{\boldsymbol{\pi}}^{\star}-{\boldsymbol{\pi}}^{\prime}_{2})-a{\boldsymbol{\lambda}}^{\top}({\boldsymbol{\pi}}^{\prime}_{1}-{\boldsymbol{\pi}}^{\prime}_{2})\leq c$. Therefore, the set $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$ is also bounded and convex in $\boldsymbol{\pi}$}.
Also, since we are working with optimistic estimate of $A$, the set $\hat{\mathcal{F}}$ will always be non-empty, because we will find at least one $\tilde{\mathbf{A}}_{0}$ which is non-singular and it’s inverse exists.
@@
    $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$ is also bounded and convex in $\boldsymbol{\pi}$}.
@@
$\text{where, }\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{A}^{i}\hat{\boldsymbol{\pi}}^{\star}\}$
Where, $\hat{\boldsymbol{\pi}}^{\star}$ is the $r$-optimal feasible policy.
[…]
@@
$\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{A^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}{A^{\prime}}\boldsymbol{\omega}$
$=$ $\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\,.$
@@ Appendix D Lagrangian Relaxation of Projection Lemma: Proof of Theorem 2
For a sequence $\{\hat{\mathcal{F}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{{\boldsymbol{\lambda}}}_{t}\}_{t\in\mathbb{N}}$, we first show that (a) $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, (b) $\boldsymbol{\lambda}$$^{\star}$ is unique, and (c) $\lim_{t\rightarrow\infty}\hat{{\boldsymbol{\lambda}}}_{t}\rightarrow{\boldsymbol{\lambda}}^{\star}$. Thus, for any $\boldsymbol{\omega}\in\mathcal{F}$ and $\boldsymbol{\mu}$},
@@
where ${\boldsymbol{\lambda}}^{*}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$.
@@
Statement (a): Convergence of the limit $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}$. To begin with the proof of the first statement of Theorem 2 we leverage the results stated in Theorem 10. Let $H(\tilde{A})\mathrel{\triangleq}\left\{{\boldsymbol{\pi}}\in\Delta_{K}:\tilde{A}{\boldsymbol{\pi}}\leq 0\right\}$ and the set function $\tilde{A}\rightarrow H(\tilde{A})\cap C$ where $C=\hat{\mathcal{F}}$ is a non-empty compact (proven in Section C subset of $\Delta_{K}$ Then the set $H(\tilde{A})\cap C$ can be written as
$H(\tilde{A})\cap C=\left\{{\boldsymbol{\pi}}\in\hat{\mathcal{F}}:\tilde{A}{\boldsymbol{\pi}}\leq 0\right\}$
To apply Theorem 10, $\{\tilde{A}^{r},r\in\mathbb{N}\}$, must be a convergent sequence of affine function. It is evident that $\tilde{A}^{r}$ for any $r\in\mathbb{N}$ is an affine function since $A$ is linear in $A$ and the induced pessimism works as a translation. Then we can proceed to the next part of the proof of statement 1 where we prove that $\{\tilde{A}^{r}\}_{r\in\mathbb{N}}$ is a convergent sequence of functions. For ease of notation we will denote $\tilde{A}_{t}$ for the t-th element of the sequence $\{\tilde{A}^{r}\}_{r\in\mathbb{N}}$ for $t\in\mathbb{N}$.
@@
[…]
where $Z\mathrel{\triangleq}\frac{\hat{A}_{t}^{i}-A^{i}}{\sigma(\hat{A}_{t}^{i})}$ and $\Phi(.)$ is the cumulative distribution function of $\mathcal{N}(0,1)$ distribution. $\lim_{t\rightarrow\infty}f(t,\delta)\|\boldsymbol{\omega}_{t}\|_{\Sigma_{t}^{-1}}\rightarrow 0$ since $\sigma(\hat{A}_{t}^{i})=\mathcal{O}(\sqrt{\frac{\log t}{t}})$. Leveraging CLT at this point we say
$\hat{A}_{t}^{i}\xrightarrow{d}A^{i},\forall i\in[d]$
[…]
It implies that $\{\tilde{A}_{r}\}_{r\in\mathbb{N}}$ is a convergent sequence of function for $A$. Now, we use Theorem 10 and get the following properties of the feasible set.
@@
    $H(\tilde{A})\cap C\subset\lim_{r\rightarrow\infty}H(\tilde{A}^{r})\cap C$
@@
    $\lim_{r\rightarrow\infty}H(\tilde{A}^{r})\cap C$ is a closed convex superset of $H(\tilde{A})\cap C$.
@@
    $H(\tilde{A})\cap C$ has non-empty interior because of the feasibility condition and no component in $\tilde{A}$ is identically $\mathbf{0}$.
    $\lim_{r\rightarrow\infty}H(\tilde{A}^{r})\cap C=H(\tilde{A})\cap C$